OpenStudy (anonymous):

HELP 2 variable: the length of a red rectangle is 4 m more than its width. A blue rectangle, which is twice as wide and 2 m longer than the red one, has a perimeter of 42 m. Find the dimensions of both rectangles

OpenStudy (anonymous):

@jim_thompson5910 hi again. um i need help D:

jimthompson5910 (jim_thompson5910):

"the length of a red rectangle is 4 m more than its width" L = W+4

jimthompson5910 (jim_thompson5910):

"A blue rectangle, which is twice as wide and 2 m longer than the red one" blue rectangle has a width twice as wide width of blue rectangle = 2W and 2 m longer length of blue rectangle = L + 2

jimthompson5910 (jim_thompson5910):

blue rectangle has "perimeter of 42 m" P = 2*(length + width) P = 2*(L + 2 + 2W) P = 2*(W+4 + 2 + 2W) ... plug in L = W+4 42 = 2*(W+4 + 2 + 2W) solve for W